OpenStudy (anonymous):

The difference between the square of one even number and the square of the next smaller even number is 52. What is the larger even number? Is it x-x^2=52 am I in the right track?

OpenStudy (anonymous):

no, the way to approach this problem is to set x = a number (odd/even doesnt matter) we know for sure, 2x = an even number so now we have the problem (2x)^2 - (2x-2)^2 = 52 solve for x, then plug back into 2x and 2x-2 for the even numbers, hope this helped!

hero (hero):

x^2 - (x - 2)^2 = 52 You don't need to use 2x to solve this

hero (hero):

x^2 - ((x-2)(x-2)) = 52 x^2 -(x(x-2)-2(x-2)) = 52 x^2 -(x^2 -2x -2x + 4) = 52 x^2 - x^2 +4x - 4 = 52 0 + 4x = 52 + 4 4x = 56 x = 14 x - 2 = 12 14^2 - (14 - 2)^2 = 52 14^2 - (12)^2 = 52 14^2 - 12^2 = 52 196 - 144 = 52 52 = 52 Therefore the consecutive numbers are 14 and 12

OpenStudy (anonymous):

@Hero thank so much for helping me out!